OpenStudy (anonymous):

Precalc: can someone help me find the directrix and focus of this parabola and then graph it? I'm not sure how to do it x^2-6y=0

OpenStudy (anonymous):

Use mathqu homework support app where you get maths answers step by step info from live tutor search for mathqu in App Store

OpenStudy (anonymous):

\[x^2=6y\] it is an upward parabole directrix \[y=-\frac{ 6 }{ 4 }=-\frac{ 3 }{ 2 },~or~2y+3=0\] focus is \[\left( 0,\frac{ 3 }{ 2 } \right)\]